# Evaluating AI systems under uncertain ground truth: a case study in dermatology

### v1

@@ 2.2 Statistical model
The statistical model $p(\lambda|b,x)$ informally introduced above and summarized in Figure 4 is at the core of our framework. Essentially, we propose to replace deterministic aggregation of annotator opinions with a statistical model. To this end, we view aggregation as computing the posterior distribution $p(\lambda|b,x)$ over plausibilities, given annotations $b$ and observations $x$. In all the examples discussed within this paper, we make the simplifying assumption $p(\lambda|b,x)=p(\lambda|b)$ such that plausibilities are inferred solely from the annotator opinions. Then, plausibilities represent distributions over labels:
@@
We compute this posterior by specifying the annotation process, i.e., $p(b|\lambda)$ and assuming a prior $p(\lambda)$ independent of the input $x$. The annotation process specifies how we expect experts to provide their annotations if we knew the underlying distribution $\lambda$ over labels. Additionally, we assume a reliability parameter as part of our statistical model. In practice, this often corresponds to a temperature parameter in $p(b|\lambda)$ (e.g., $\gamma$ in our toy example from Figure 2). However, we interpret it as quantifying the prior trust we put in the annotators. As fixing this parameter based on domain expertise or data is challenging, we treat it as a free parameter that is to be explored during evaluation.
On CIFAR10, we assumed each annotator to provide the top-1 label. Thus, it was simple to derive the posterior in closed form for a Dirichlet prior distribution. Here, the plausibilities $\lambda$ explicitly correspond to the categorical distribution over classes $p(y|x)$. In other cases, where annotations do not directly match the label space (see Appendix A for examples), this statistical model might be more complex. For our case study, a skin-condition classification problem detailed in Section 3, the expert annotations are partial rankings instead of top-1 labels. In this case, the plausibilities $\lambda$ approximate the categorical distribution $p(y|x)$ as for CIFAR10 but we will rely on a Plackett–Luce model for $p(b|\lambda)$ (Plackett, 1975; Luce, 2012). Unfortunately, this means that the posterior $p(\lambda|b)$ is not available in closed-form so specific techniques are needed to sample from it; see Appendix B.
@@
Figure 4: Graphical model corresponding to the statistical aggregation model at the core of our framework: This describes the joint distribution of the observation $x$, plausibilities $\lambda$, annotations $b=(b^{1},...,b^{R})$ and label $y$. We observe $x$ and $b$ from which one can infer plausibilities $p(\lambda|b,x)$ and label $p(y|\lambda)$, see text. As discussed in the text, in order to simplify inference of $\lambda$, we assume a simplified model with $p(\lambda|b,x)=p(\lambda|b)$.
@@
Our model in Figure 4 comes with assumptions and limitations that are important to highlight. First, we assume the annotators to be conditionally independent given the plausibilities. This is a simplification and there is significant work in crowd sourcing and truth discovery considering alternative models (Yan et al., 2014). Second, as discussed previously, in all our examples, we consider a simplified version of the model in Figure 4, assuming conditional independence $p(\lambda|b,x)=p(\lambda|b)$. This makes inferring the posterior over $\lambda$ easier but clearly reduces our ability to model input dependent uncertainty and thereby disentangle the different sources of uncertainty outlined above. This means that $p(\lambda|b)$ can only capture the uncertainty present in the annotations. This reduces our ability to decide whether the uncertainty stems mainly from limited observational information or from the annotation process. For example, it is difficult to distinguish between an inherently ambiguous example where we observed low disagreement by chance and an actually unambiguous example with low disagreement. However, this distinction would be extremely valuable, e.g., to inform relabeling. Finally, it is important to realize that this introduces a model assumption in our certainty and performance metrics. For example, there is no guarantee that $p(y|b,x)$ converges to the true $p(y|x)$ as the number of annotators goes to infinity as the annotation model can be mis-specified. However, all existing benchmarks are based on an assumed annotation model. That this assumption is implicit and often unacknowledged does not change the fact that it has tangible effects on evaluation.
@@ 2.3 Annotation certainty
To formalize our measure for annotation uncertainty as described above, we consider a fixed but arbitrary label $y\in[K]$. This could be a deterministically aggregated ground truth as on CIFAR10 or, in Section 2.4, a prediction from a classifier. Then, we informally define the certainty for a specific label $y$ as the probability that $y$ corresponds to the top-1 label of the plausibilities $\lambda$, given the input $x$ and annotations $b$ as well as the chosen statistical model. More formally, we can write this as an expectation over $p(\lambda|b,x)$:
$\text{Certainty}(y;b,x)=p(y=\arg\max_{j}\lambda_{j})=\mathbb{E}_{p(\lambda|b,x)}\Big{\{}\delta\left[{y=\arg\max_{j}\lambda_{j}}\right]\Big{\}}.$ (2)
where $\delta[\cdot]\in\{0,1\}$ is the indicator function for an event. In practice, we compute it using a Monte Carlo average
@@
This estimates certainty for a specified label $y$. To summarize annotation certainty for an example, we compute the maximum certainty across all possible labels $y$:
@@
Here, the subscript $1$ indicates that annotation certainty refers to the top-1 label from plausibilities. However, we can similarly generalize this to sets of labels $C$, e.g., top-$j$ sets with $|C|=j$:
@@
with $\text{Certainty}(C;b,x)=\mathbb{E}_{p(\lambda|b,x)}\Big{\{}\delta\left[{C=Y_{\text{top}-|C|}(\lambda)}\right]\Big{\}}.$ (6)
This certainty measure can also be estimated using $M$ Monte Carlo samples. This prevents us from having to enumerate all $K!/(j!(K-j)!)$ subsets of size $j$ as seemingly implied by (5), which would be prohibitive. Instead, we only need to consider top-$j$ sets corresponding to the $M$ samples. However, larger $j$ might require a higher $M$ for reliable estimates of annotation certainty. To measure annotation certainty of a dataset, we can average AnnotationCertainty($b,x$) across examples.
There are some caveats of our measure of annotation certainty to be aware of, most of them depending on the aggregation procedure $p(\lambda|b,x)$. For example, Equation (5) is always $1$ for any input where the model $p(\lambda|b,x)$ is a point mass; i.e., it gives a single deterministic point estimate $\hat{\lambda}$ (subject to there being ties, which we ignore for simplicity). This is, by construction the case in any deterministic aggregation procedure for human annotations. Also, with an infinite number of annotators $p(\lambda|b,x)$ will typically converge to a point mass $\hat{\lambda}$ such that annotation certainties converge to $1$. However, this does not mean that the plausibilities correspond to an unambiguous one-hot distribution over classes. Finally, annotation certainties for the same problem depend on the used statistical aggregation process and the reliability we attribute to annotators. This also means that annotation certainty is not a metric that is to be minimized. This is trivially possible, for example, by putting infinite trust in annotators and working with a point estimate $\hat{\lambda}$ (infinite reliability $\gamma$ in Figure 2).
Considering the 3-simplex in Figure 2, annotation certainty measures the impact of the spread in plausibilities on the top-$j$ labels. Thereby, annotation certainty implicitly also captures inherent uncertainty: even a small spread in plausibilities changes the top-$j$ labels more often for inherently uncertain examples, i.e., plausibilities close to the center of the simplex. More generally, this means that annotation certainty will be high for easy, unambiguous examples that all annotators agree on. The reverse, however, is not necessarily true: low annotation certainty can indicate high inherent uncertainty but it does not necessarily have to. For example, many experienced annotators consistently disagreeing might indicate high inherent uncertainty; however, annotators might also disagree for other reasons such as unclear annotation instructions, even if the example is generally easy to classify. This also explains our naming: annotation certainty explicitly measures annotation uncertainty and only implicitly accounts for inherent uncertainty. Partly, this also stems from our assumption that $p(\lambda|b,x)=p(\lambda|b)$ in Section 2.2, making it more difficult to disentangle annotation and inherent uncertainty.

### v2

@@ 2.2 Statistical model
The statistical model $p(\lambda|b,x)$ informally introduced above and summarized in Figure 4 is at the core of our framework. Essentially, we propose to replace deterministic aggregation of annotator opinions with a statistical model. To this end, we view aggregation as computing the posterior distribution $p(\lambda|b,x)$ over plausibilities, given annotations $b$ and observations $x$. Then, plausibilities represent distributions over labels:
@@
In all the examples discussed within this paper, we make the simplifying assumption $p(\lambda|b,x)=p(\lambda|b)$ such that plausibilities depend only from the annotator opinions. Computing the posterior distribution $p(\lambda|b)$ is a central task for our purposes. We compute this posterior by specifying the annotation process, i.e., $p(b|\lambda)$ and assuming a prior $p(\lambda)$ independent of the input $x$. The annotation process specifies how we expect experts to provide their annotations if we knew the underlying distribution $\lambda$ over labels. Additionally, we assume a reliability parameter as part of our statistical model. In practice, this often corresponds to a temperature parameter in $p(b|\lambda)$ (e.g., $\gamma$ in our toy example from Figure 2). However, we interpret it as quantifying the prior trust we put in the annotators. As fixing this parameter based on domain expertise or data is challenging, we treat it as a free parameter that is to be explored during evaluation.
In the toy example of Section 2.1, we assumed that each annotator provides the top-1 label. Thus, it was simple to derive the posterior in closed form for a Dirichlet prior distribution. Here, the plausibilities $\lambda$ explicitly correspond to the categorical distribution over classes $p(y|x)$. In other cases, where annotations do not directly match the label space (see A for examples), this statistical model might be more complex. For our case study, a skin-condition classification problem detailed in Section 3, the expert annotations are partial rankings instead of top-1 labels. In this case, the plausibilities $\lambda$ still approximate the categorical distribution $p(y|x)$ but we need to rely on more complex models for $p(b|\lambda)$. This might entail that that the posterior $p(\lambda|b)$ is not available in closed-form, requiring specialized techniques for sampling.
@@
Fig. 4: Graphical model corresponding to the statistical aggregation model at the core of our framework: This describes the joint distribution of the observation $x$, plausibilities $\lambda$, annotations $b=(b^{1},...,b^{R})$ and label $y$. We observe $x$ and $b$ from which one can infer plausibilities $p(\lambda|b,x)$ and label $p(y|\lambda)$, see text. As discussed in the text, in order to simplify inference of $\lambda$, we assume a simplified model with $p(\lambda|b,x)=p(\lambda|b)$.
@@
Beyond the specific choice of $p(b|\lambda)$, we assume the annotators to be conditionally independent given the plausibilities. This is a simplification and there is significant work in crowd sourcing and truth discovery considering alternative models [116]. Moreover, in our experiments, we consider a simplified version of the model in Figure 4 that assumes conditional independence $p(\lambda|b,x)=p(\lambda|b)$. This makes inferring the posterior over $\lambda$ easier but clearly reduces our ability to model input dependent uncertainty and thereby disentangle the different sources of uncertainty. For example, it is difficult to distinguish between an inherently ambiguous example where we observed low disagreement by chance and an actually unambiguous example with low disagreement. However, this distinction would be extremely valuable, e.g., to inform relabeling.
One might argue that we introduce a model assumption into evaluation and there is no guarantee that $p(y|b,x)$ converges to the true $p(y|x)$ as the number of annotators goes to infinity. However, it is important to realize that almost all existing benchmarks are based on an assumed annotation model. This is because these benchmarks obtain ground truth labels through expert annotations. That the assumed annotation model is implicit and often unacknowledged in the corresponding publications and leaderboards does not change the fact that it has tangible effects on evaluation. We make this model assumption explicit and actually use it for our advantage: quantifying the underlying ground truth uncertainty.
2.3 Model for differential diagnosis annotations
To apply this framework to dermatology, we model differential diagnoses as partial rankings and discuss meaningful aggregation models $p(\lambda|b,x)$ that facilitate sampling plausibilities. Specifically, we follow [58], and use inverse rank normalization (IRN) as a deterministic aggregation baseline. We then propose a probabilistic variant of IRN as a first, simple statistical aggregation model before considering a more sophisticated Plackett–Luce model [75, 61].
For each case, each annotator selects a subset of conditions they think might be the true diagnosis. In addition, they assign a confidence score to each condition, indicating how likely they think that condition is the true diagnosis. The subsets of conditions selected by different annotators may differ in size. Moreover, the confidence scores are usually not comparable across annotators which is why they are instead used to produce a ranking of the conditions for each annotator [58]. This results in partial rankings because there may be ties among conditions (i.e., conditions with the same confidence score) and the large majority of possible conditions remains unranked. Formally, we write
$b=(b_{{}{}{1}}\succ b_{{}{}{2}}\ldots\succ b_{{}{}{L}}).$ (2)
We call groups of tied conditions “blocks”, and each block contains one or more labels, $b_{{}{}{i}}=\{{\sigma_{{i}{1}}},\ldots,{\sigma_{{i}{|b_{{}{}{i}}|}}}\}$ with ${\sigma_{{i}{j}}}\in[K]$ and $K$ being the number of classes. Here, $[P]:=\{1,...,P\}$ for integer $P$ and $b_{{}{}{i}}\succ b_{{}{}{j}}$ indicates that classes in the $i$-th block are ranked higher than those in the $j$-th block (according to the confidence scores). We assume the blocks are non-empty, mutually exclusive, i.e., $\emptyset\neq b_{{}{}{i}}\subseteq[K]$ and $b_{{}{}{i}}\cap b_{{}{}{j}}=\emptyset$, and that $\bigcup_{i=1}^{L}b_{i}=[K]$. Note that there might exist multiple equivalent permutations within the individual blocks $b_{i}$ such that we use $\mathcal{S}(b)$ to denote the set of all permutations compatible with Equation (2). In the literature discussing ranking metrics and aggregation [86, 111, 94, 114], partial rankings are also commonly referred to as top-$k$ rankings with ties. While some metrics such as Kendall’s tau and Spearman’s footrule have been adapted to ties [33, 106], modeling partial rankings is generally non-trivial. This also holds for any aggregation model $p(\lambda|b,x)$ as we will see in Section 2.3.3.
$b=($ $b_{1}=\{\textit{Hemangioma}\},$
[…]
We interpret the above example as follows: the annotator declares that Hemangioma is their first choice, followed by either Melanocytic Nevus, Melanoma, or O/E. However, for the latter three conditions, the annotator is unable to say which one is more likely than the others. That is, these three conditions are tied for the second rank. Implicitly, this means that all $3!=6$ possible permutations of the second rank would be equivalent. Moreover, annotators usually ignore a large set of conditions. Strictly speaking, the ignored conditions can be thought of as another block
$b_{3}=\{\text{any other condition}\}$
with the last block $b_{L}=b_{3}$ capturing all unranked classes.
2.3.1 Inverse rank normalized (IRN) aggregation
A heuristic method for aggregating differential diagnoses to arrive at a plausibility point estimate is Inverse Rank Normalization (IRN) [58]. Given a partial ranking $b$ from a single annotator with $L$ blocks as in Equation (2), we define the unnormalized IRN score of a condition $y$ as
$\overline{\text{IRN}}(y;b)=\sum_{i=1}^{L-1}\frac{\nicefrac{{1}}{{i}}}{|b_{{}{i}{}}|}\delta[y\in b_{{}{i}{}}].$ (3)
where $\delta$ is an indicator function where $\delta[S]$ is one when the statement $S$ is true and zero otherwise. Unnormalized IRN assigns a weight $\nicefrac{{1}}{{i}}$ to each block $b_{{}{i}{}}$ for $i\in[L-1]$ and a weight of 00 for all the classes in the last block $b_{{}{}{L}}$ of unranked classes. The weight is distributed equally across all classes in a block, i.e., class ${\sigma_{{i}{j}}}$ obtains weight $\nicefrac{{{1/i}}}{{|b_{{}{i}{}}|}}$. Given $R$ partial rankings $b^{1},\ldots,b^{R}$ from $R$ annotators, we define the unnormalized IRN score of a condition $y$ as
$\overline{\text{IRN}}(y;b^{1},\ldots,b^{R})=\sum_{r=1}^{R}\overline{\text{IRN}}(y;b^{r})$ (4)
$\text{IRN}(y;b^{1},\ldots,b^{R})=\frac{\overline{\text{IRN}}(y;b^{1},\ldots,b^{R})}{\sum_{k=1}^{K}\overline{\text{IRN}}(k;b^{1},\ldots,b^{R})}.$ (5)
Note that normalization comes after aggregation rather than normalizing per annotator $r$ first and then averaging.
Overall, IRN can be viewed as an adhoc point estimate of the plausibilities $\lambda$ without an explicit generative model for $p(\lambda|b,x)$. This also implies that we cannot sample plausibilities from $p(\lambda|b,x)$. Moreover, because it is deterministic, it also has to resolve ties deterministically. However, distributing a rank’s weight equally across tied conditions can lead to unintuitive plausibilities (see Figure 5 for an example and discussion). In contrast, the next section describes a probabilistic version of IRN that allows to sample plausibilities.
2.3.2 Probabilistic IRN (PrIRN) aggregation
To derive a probabilistic variant of IRN (PrIRN), we interpret the calculation of the unnormalized IRN as a counting method for accumulating the statistics of a histogram on discrete data. Then, the normalized $\text{IRN}(y;b^{1},\ldots,b^{R})$ corresponds to a maximum likelihood estimate of the parameters of a multinomial distribution (the sampling distribution of a histogram) given the accumulated statistics. In this spirit, we define the sampling distribution $p(\lambda|b^{1},\ldots,b^{R})$ directly as a Dirichlet distribution $\mathcal{D}(\gamma\hat{\lambda})$ where $\hat{\lambda}_{y}=\text{IRN}(y;b^{1},\ldots,b^{R})$ and $\gamma>0$ (the concentration parameter) represents our reliability parameter. Here, the $\mathcal{D}(\gamma\hat{\lambda})$ distribution is the true posterior distribution of the parameters of a categorical distribution if labels were drawn i.i.d. and the observed counts would be $\gamma\hat{\lambda}$. Then, the reliability parameter $\gamma$ can be interpreted as the number of annotators: with increasing number of annotators we have less annotation uncertainty and for $\gamma\rightarrow\infty$ the Dirichlet distribution degenerates to a point mass on $\hat{\lambda}$ – the IRN estimate.
2.3.3 Plackett–Luce (PL) aggregation
One conceptual problem of PrIRN is that its weighting and aggregation choices appear to be fairly arbitrary and the corresponding generative model for the annotation process based on a categorical distribution is not realistic for differential diagnoses. An alternative is defining a generative model, that takes the ordering into account: when raters provide a partially ordered list of conditions, they exclude already mentioned conditions when choosing the next option.
The Plackett–Luce (PL) [75, 61] model is a natural choice for this setting. The PL distribution models how annotators, given the true plausibilities $\lambda$, draw full rankings by sampling without replacement from all conditions. This model is very appropriate for our problem. Unfortunately, in a Bayesian framework, the resulting posterior distribution of the plausibilities cannot be derived analytically. However, efficient Gibbs sampling algorithms [15, 37] and optimization procedures [46] do exist. Here, we extend the Gibbs sampler of [15] to partial rankings.
In the following, we assume a single case for brevity, with the annotator being indexed by $r\in[R]$. The inferential procedure we describe can be repeated for any other case independently. We also assume that
$p({\sigma}|\lambda)=\prod_{r=1}^{R}p({\sigma^{{r}}}|\lambda),$ (6)
meaning that the annotators’ rankings are independent, conditioned on plausibilities. Then, moving to the description of $p({\sigma^{{r}}}|\lambda)$, the classical PL model makes the assumption that each annotator $r$ reports a full ranking of classes by sampling without replacement, proportional to the plausibilities $\lambda$. So the first class ${\sigma^{{r}}_{{}{1}}}$ is drawn with probability
$p({\sigma^{{r}}_{{}{1}}}=k)=\frac{\lambda_{k}}{Z}$
$\mathcal{L}_{r}(\lambda)=\log\lambda_{{\sigma^{{r}}_{{}{1}}}}-\log Z+$ $\log\lambda_{{\sigma^{{r}}_{{}{2}}}}-\log(Z-\lambda_{{\sigma^{{r}}_{{}{1}}}})+\dots$
$+\log\lambda_{{\sigma^{{r}}_{{}{K}}}}-\log\left(Z-\sum_{k=1}^{K-1}\lambda_{{\sigma^{{r}}_{{}{k}}}}\right).$
The inferential goal of interest is estimating the latent plausibilities $\lambda$ based on the annotations $b$ in the form of full rankings $\sigma$. This can be achieved by maximizing the likelihood $\mathcal{L}(\lambda)$, or in a Bayesian way by assigning a prior distribution $p(\lambda)$ and sampling from the posterior distribution $p(\lambda|{\sigma})\propto p(\lambda)\prod_{r}p({\sigma^{{r}}}|\lambda)$. As the maximum likelihood solution would provide a point estimate of the plausibilities, similar to IRN, here we focus on the sampling approach.
We introduce the equivalent of a reliability parameter $\gamma$ for the PL distribution by defining a new tempered PL distribution with parameters $\bar{\lambda}=\lambda^{\gamma}$ and normalization $\bar{Z}=\sum_{k}\bar{\lambda}_{k}=\sum_{k}\lambda^{\gamma}_{k}$. Unfortunately, the resulting tempered model no longer enjoys the conjugacy properties needed for deriving a Gibbs sampling algorithm. While it is possible to use more generic sampling methods such as Hamiltonian Monte Carlo (HMC) for this tempered PL model, qualitatively we obtain the same concentration effect by simply repeating the partial rankings for each annotator such that the likelihood contribution of each annotator to the posterior is $\gamma\mathcal{L}_{r}(\lambda)$ where $\gamma$ is a small integer. This also matches the intuition of the reliability parameter $\gamma$ in PrIRN from the previous section.
The PL model as introduced above defines a distribution over full rankings. While it is straightforward to compute the probability of a given full ranking under the PL distribution, partial rankings – especially with larger block sizes – provide a computational challenge as we need to consider all the permutations compatible with the partial rankings. This would amount to a computational complexity that scales factorially in the block size. In contrast, in B.1, we describe a procedure for computing the exact likelihood for partial rankings that scales only as $O(2^{|B|})$ with the block sizes $|B|$ making it practical for typical differential diagnoses.
@@ 2.4 Annotation certainty
Having formalized two aggregation models $p(\lambda|b)$ to use for differential diagnosis annotations, we now formalize our measure of annotation uncertainty. We first consider a label $y\in[K]$, and define the certainty for a specific label $y$ as the probability that $y$ corresponds to the top-1 label of the plausibilities $\lambda$, given the input $x$ and annotations $b$ under the chosen statistical model for aggregation. More formally, we can write this as an expectation under $p(\lambda|b,x)$:
$\text{Certainty}(y;b,x)$ $=p(y=\arg\max_{j}\lambda_{j})$ (7)
$=\mathbb{E}_{p(\lambda|b,x)}\Big{\{}\delta\left[{y=\arg\max_{j}\lambda_{j}}\right]\Big{\}}.$
In practice, we compute this expectation using a Monte Carlo average
$\text{Certainty}(y;b,x)\approx\frac{1}{M}\sum_{m=1}^{M}$ $\delta\left[{y=\arg\max_{j}\lambda_{j}^{(m)}}\right],$ (8)
@@
This estimates certainty for a specified label $y$ for a specific example. To summarize the overall annotation certainty for the example, we compute the maximum certainty across all possible labels $y$:
@@
Here, the subscript $1$ indicates that annotation certainty refers to the top-1 label from plausibilities.
An important feature of the above definition is that we can generalize it to sets of labels $C$, e.g., top-$j$ sets with $|C|=j$:
@@
$\text{with}\quad\text{Certainty}(C;b,x)=\mathbb{E}_{p(\lambda|b,x)}\Big{\{}\delta\left[{C=Y_{\text{top}-|C|}(\lambda)}\right]\Big{\}},$ (11)
where $Y_{\text{top}-|C|}(\lambda)$ refers to the set that includes $C$ most plausible labels based on $\lambda$. This certainty measure can also be estimated using $M$ Monte Carlo samples. This prevents us from having to enumerate all $K!/(j!(K-j)!)$ subsets of size $j$ as seemingly implied by (10), which would be prohibitive. Instead, we only need to consider top-$j$ sets corresponding to the $M$ samples. However, larger $j$ might require a higher $M$ for reliable estimates of annotation certainty. To measure annotation certainty of a dataset as a whole, we can average AnnotationCertainty($b,x$) across examples.
Note that this measure depends on the aggregation procedure $p(\lambda|b,x)$. For example, Equation (10) is always $1$ for any input where the model $p(\lambda|b,x)$ is a point mass; i.e., it gives a single deterministic point estimate $\hat{\lambda}$ (subject to there being ties, which we ignore for simplicity). This is, by construction the case in any deterministic aggregation procedure for human annotations.
To understand intuitively what the measures are calculating, considering the simplex in Figure 2. Each point on this simplex corresponds to an individual plausibility vector. For a given case, if we use a deterministic aggregation methods, the plausibility vector is a single point estimate and the top-$1$ label is associated with one of the three sectors shown on the simplex. The annotation certainty measures the impact of the spread in plausibilities on the top-$j$ labels. Thereby, annotation certainty implicitly also captures inherent uncertainty: even a small spread in plausibilities changes the top-$j$ labels more often for inherently uncertain examples, i.e., plausibilities close to the center of the simplex. More generally, this means that annotation certainty will be high for easy, unambiguous examples that all annotators agree on. The reverse, however, is not necessarily true: low annotation certainty can indicate high inherent uncertainty but it does not necessarily have to. For example, many experienced annotators consistently disagreeing might indicate high inherent uncertainty; however, annotators might also disagree for other reasons such as unclear annotation instructions, even if the example is generally easy to classify. This also explains our naming: annotation certainty explicitly measures annotation uncertainty and only implicitly accounts for inherent uncertainty. Partly, this also stems from our assumption that $p(\lambda|b,x)=p(\lambda|b)$ in Section 2.2, making it more difficult to disentangle annotation and inherent uncertainty.